# Explaining Time Series Classifiers with PHAR: Rule Extraction and Fusion from Post-hoc Attributions

### v3

@@ 3.1 Overview of the PHAR Framework
The PHAR (Post-hoc Attribution Rules for time series) framework addresses the demand for semantically meaningful, stable, and cognitively efficient explanations in time series classification. In PHAR, we transform post-hoc, instance-level numeric feature attributions into concise, human-readable rule-based explanations and integrate outputs from multiple explainers enabling assessment of the quality of explanations and their visualization. First, PHAR selects the most salient temporal intervals—and, in multivariate settings, the most informative channels—so that rules focus on key patterns. Next, these intervals are translated into interpretable “IF–THEN” rules that preserve the model’s predicted class. An optional fusion step then merges rules from obtained from numeric attributions, Anchor explanations [29], and potentially expert rules, optimizing for coverage, confidence, and simplicity. Fusing explanations via techniques such as intersection, union, and weighted aggregation ensures faithfulness to DL classifier predictions and usability for end-users. Finally, to reduce extraneous cognitive load, PHAR visualizes each rule with semi-factual plots that contrast actual and hypothetical behaviors, highlighting how small changes would affect the model’s decision. Our framework is structured around 4 Research Questions that guide the transformation, fusion, and visualization of rule-based explanations for time series classification.
@@
RQ3: How can fusion methods combining rules from LIME, SHAP, and Anchor generate a single optimized rule per instance? Can this transformation also refine rules within a single explainer? Whether merging multiple perspectives into one rule per instance improves coverage, confidence, and explanation ratio without excessive complexity?
RQ4: How can visualizations support the interpretability of rule-based explanations in TS models? Rules may span multiple time steps and features (in multivariate TS). We explore whether overlaying rule-based explanations on TS plots helps illustrate their structure.
Figure 2: Proposed framework with highlighted research questions RQ1…RQ4.
Figure 2 illustrates the PHAR framework with those Research Questions. A deep classifier first assigns each input series: univariate or multivariate, to a class. Post-hoc explainers SHAP and LIME yield numerical attributions that we convert into interval rules of the form $r:\;x_{c}^{(t)}\in(l_{c}^{(t)},h_{c}^{(t)}]\;\Rightarrow\;\hat{y},\quad\text{conf}(r)=\gamma,\;\text{cov}(r)=\kappa,$ where $x_{c}^{(t)}$ denotes "channel" $c$ (in case of multivariate TS) at timestep $t$, $(l_{c}^{(t)},h_{c}^{(t)}]$ is the selected interval, and $\gamma,\kappa$ are the rule’s confidence and coverage scores. This transformation addresses RQ1 (Section 3.3). In parallel, Anchor generates rule-based explanations directly. We evaluate all candidate rules using metrics such as coverage, confidence, and feature count to address RQ2 (Section 3.3.1). Then we apply multiple fusion strategies (Intersection, Union, Weighted, Lasso, Lasso global, and Best), controlled by these metrics, to select a single, optimized rule per instance (RQ3, Section 3.4). Finally, to support RQ4, we visualize each rule as vertical intervals overlaid on the original TS, highlighting feature–timestep relationships (Section 4.7 and Appendix A.4).
@@ 3.2 Rule-based explanations
In Davis et al. [10], rules are expressed in the "IF–THEN" form augmented with a certainty factor. The premise is a conjunction of predicate evaluations $p_{j}(o_{j},a_{j})=v_{j}$ for $j=1,\dots,n$, and the conclusion $q(o^{\prime},a^{\prime})=v^{\prime}_{\mathrm{out}}$ is drawn with a strength given by multiplying the aggregated premise confidence by $\mathrm{CF}_{i}$:
@@
Here, $p_{j}$ denotes a predicate function applied to object $o_{j}$ and attribute $a_{j}$, $v_{j}$ is its value, and the certainty factor $\mathrm{CF}_{i}$ quantifies the inference reliability. Building on this idea, we define our PHAR expert-system rule format format as in [16, 24], where each rule $R_{i}$ comprises a set of logical conditions $W_{1},\dots,W_{n}$, a conclusion $W_{\mathrm{out}}$, and an inference semaphore $S_{i}$ representing the combined confidence $c_{i}$ obtained with the method described in Section 3.3.1:
@@
In this schema, each $W_{j}$ is a logical condition derived from feature attributions, $W_{\mathrm{out}}$ is the predicted class or decision, and $S_{i}=c_{i}$ signals the rule’s confidence.
Our approach follows established rule classifications in knowledge-based systems [22]. PHAR rules are best described as derivation and classification rules. Derivation rules define logical conditions under which a conclusion is inferred from given premises, in our case TS feature conditions leading to model predictions. Classification rules assign instances to classes based on these conditions. From the perspective of logical inference, our rules are probabilistic deductive rules. Deductive rules infer conclusions from premises. If premises hold, the conclusion follows deductively. Given a feature set and its constraints, our rules classify instances under a probabilistic confidence threshold rather than absolute certainty. Unlike classical deduction, our rules incorporate confidence, applying deterministically within probability constraints. Furthermore, our rules belong to attributive rule languages [22]. These languages encode knowledge via attributes and values, common in decision tables, rule-based systems, and databases. They enable intuitive representation and efficient manipulation. For example, the rule if $\mathrm{temperature}(t=10)>50$ and $\mathrm{pressure}(t=11)<10$, then class = failure [$c=0.85$], where $t=11$ denotes the timestep index, provides a human-readable decision boundary that replaces raw numeric attribution scores.
Converting scores from explainers which produce raw post-hoc numerical attributions, such as SHAP and LIME, into a structured rule format makes it possible to compare and merge rules from different explainers, resolve conflicts when features are correlated, and engage Domain Experts in validating and refining the symbolic rules. Our approach employs rule fusion to aggregate rule sets produced by post-hoc explainers and additionally allows to incorporate native rule-based explainers like Anchor or even rules developed directly by Domain Experts. Inspired by ensemble methods [11], we call this process fusion because it goes beyond classical bagging or boosting. Fusion refines and unifies rules by weighting them according to coverage, precision, and simplicity, and by merging overlapping or redundant conditions. The resulting fused rules conform to the same expert-system format, allowing them to be easily evaluated, adjusted, visualized, and aligned with expert knowledge to support robust domain reasoning.
@@
Here, each pair $t_{j_{m}}c_{j_{m}}$ denotes “channel” $c_{j_{m}}$ (in multivariate series) at timestep $t_{j_{m}}$. The indices $j_{1},\dots,j_{k}$ are selected arbitrarily from the full set of $T$ time steps and $C$ "channels", choosing the $k+1$ pairs with the greatest impact on model predictions. Each interval $(\ell_{j_{m}},u_{j_{m}}]$ has lower bound $\ell_{j_{m}}$ and upper bound $u_{j_{m}}$. $\mathrm{CONF}(R_{i})$ is the proportion of instances satisfying the rule that are correctly classified as class $A"$, and $\mathrm{COV}(R_{i})$ is the fraction of all instances that satisfy the rule. The formal derivation of rules and their CONF and COV measures is presented in Sections 3.3 and 3.3.1, while Section 4.7 provides real-world examples of rules in this notation for selected dataset.
@@ 3.3 Conversion of numeric importance values to rules
We consider a TS dataset with $N$ samples, univariate or multivariate, and a model generating multi-class predictions. In the univariate case, each feature $f$ is indexed only by timestep, e.g. $\mathrm{t}_{i}$, whereas in the multivariate case, features are indexed by both timestep and channel, e.g. $\mathrm{t}_{i}\mathrm{c}_{v}$. Let $\{X_{n}\}$ be instances with predicted classes $y_{n}$. Predicted classes serve as ground truth, as we focus on explanation consistency with DL model, not model accuracy itself. Feature-based explainers (e.g., LIME, SHAP) assign numeric importance values to each variable per timestep. Let $e_{n,f}$ denote importance values for instance $n$ and feature $f$. To extract rules, we identify a threshold $T$ (percentile-based) defining important features. Thresholds are computed per feature (local) or across all features (global).
We first compute absolute explanation values for all training instances and features. Let $E$ be the multiset of absolute values: $E=\{|e_{n,f}|:n=1,\dots,N,\;f=1,\dots,F\},$ where $e_{n,f}$ is the explanation value for feature $f$ (e.g., $\text{t}_{i}\text{c}_{v}$ for multivariate or $\text{t}_{i}$ for univariate TS) at instance $n$. We then define a global threshold $T_{global}$ by selecting a percentile $p$: $T_{global}=\text{Percentile}(E,p).$ If we apply thresholds per feature, we instead compute a separate threshold $T_{f}$ for each feature $f$: $E_{f}=\{|e_{n,f}|:n=1,\dots,N\},\quad T_{f}=\text{Percentile}(E_{f},p).$ During inference on test instances, for each feature $f$, let $e_{n,f}$ be the explanation value for instance $n$:
@@
Features exceeding their threshold are important. For important features, we generate intervals preserving the model’s prediction. Stable intervals $(l_{f},u_{f}]$ are derived via controlled perturbation of each important feature and sampling around $x_{n,f}$. Let $F_{*}$ be important features for an instance, identified after thresholding. For each $f\in F_{*}$, we take the original value $x_{n,f}$ and generate perturbed samples: $x_{n,f}^{(m)}\sim\text{Uniform}(x_{n,f}-\Delta_{f},\;x_{n,f}+\Delta_{f}),$ where $\Delta_{f}$ is a perturbation range based on training explanation statistics (std. deviation × constant $\sigma_{p}$). Joint sampling over $f\in F_{*}$ captures nonlinear feature dependencies. For each perturbed sample $(x_{n,f}^{(m)})_{f\in F_{*}}$, we calculate the model’s prediction. Let $c_{ref}(n)$ be the original predicted class of instance $n$. If $c_{ref}(n)=y_{n}$, we determine $(l_{f},u_{f}]$ per feature $f$ preserving this class. Each $f\in F_{*}$ receives a stable interval: $f:x_{f}\in(l_{f},u_{f}],$ defining a rule $R_{n}$ for instance $n$:
@@ 3.3.1 Confidence and coverage of newly derived rules
After deriving the final rule $R_{n}$ for instance $n$, we evaluate its explanatory quality on the test set. Consider the test dataset $\{X_{test,k}\}_{k=1}^{N_{test}}$, where $N_{test}$ is the number of test instances. We identify the subset of test instances that satisfy the rule: $S(R_{n})=\{X_{test,k}:X_{test,k}\text{ satisfies }R_{n}\}.$ To measure coverage, we compute:
@@
This ratio indicates how many test instances fall into the region defined by the rule’s intervals. A higher $COV(R_{n})$ means the rule covers a larger portion of the test set.
To measure confidence, we focus on the consistency of predictions within $S(R_{n})$. Let: $T(R_{n})=\{X_{test,k}\in S(R_{n}):y_{model}(X_{test,k})=c_{ref}(n)\},$ i.e., the subset of covered test instances for which the model’s predicted class matches $c_{ref}(n)$. The confidence is then:
@@
If no test instance satisfies the rule ($S(R_{n})=\emptyset$), then $COV(R_{n})=0$ and $CONF(R_{n})$ is undefined, meaning no explanatory power. However, at least instance $n$ should satisfy the rule. If the rule covers test instances but does not match $c_{ref}(n)$, confidence decreases.
@@ 3.3.2 Hyperparameter optimization in rule transformation step
To ensure that our rule‐based explanations achieve a balance between fidelity and simplicity, we employ Optuna [^2] Optuna: https://optuna.org/ [2] to tune key hyperparameters of this step of our framework.
@@
  • $p$ (threshold percentile): Percentile for feature importance thresholds. Higher values prioritize top-scoring features thus results in lower quantity of important features.
@@
  • $\sigma_{p}$ (perturbation scale): Perturbation scale. Higher values widen the search for stable intervals.
  • $N_{p}$ (perturbation samples count): Number of perturbed samples for interval estimation. More samples improve precision but increase computational cost.
Optuna searches over the hyperparameters $p$, $\delta$, $\sigma_{p}$ and $N_{p}$ to maximize objective on a held‐out test set. During each trial, candidate configurations are evaluated by applying uniform sampling to estimate rule stability and decision region consistency, ensuring stable decision regions in the ML model’s input space. The best configuration is selected based on its performance on the test set, ensuring that the chosen thresholds generalize beyond the training data. Tuning these parameters balances feature selectivity, interval granularity, and computational cost. Table 1 (Appendix) summarizes hyperparameter ranges used in our study.
Specifically, we define an objective function maximizing coverage and confidence while penalizing empty or complex rules during numeric explanations to rule transformation step which is then feed to Optuna framework.
@@
Each instance $n$ starts with a complex rule from an explainer (e.g., Anchor, LIME, SHAP). Each rule defines intervals for features, e.g., timestep-channel $\text{t}_{i}\text{c}_{v}$ in multivariate TS or timestep $\text{t}_{i}$ in univariate TS, denoted $F(R_{n})$. Rules yield metrics: coverage $\operatorname{COV}(R_{n})$, confidence $\operatorname{CONF}(R_{n})$, and count of features which rule specifies $|F(R_{n})|$. The initial objective function is:
@@
If rule for instance $n$ is empty or absent, $M(n)=0$. Otherwise, penalties are applied when certain thresholds are not met.
Then we introduce thresholds $\tau_{conf}$, $\tau_{cov}$, and $\tau_{feat}$, which are chosen to balance rule quality and interpretability. $\tau_{conf}$ ensures a minimum confidence level for the rules, $\tau_{cov}$ prevents overly narrow rules by enforcing minimum coverage, and $\tau_{feat}$ limits the number of features in a rule to improve cognitive interpretability for end-users. We set $\tau_{conf}=0.5$ because scikit-learn and many binary classifiers use 0.5 as the default probability cutoff, aligning rule applicability with positive predictions above chance level and following standard model calibration practices. The coverage threshold $\tau_{cov}=0.01$ (1% support) corresponds to the minimum support filter commonly adopted in association rule mining to remove infrequent patterns and ensure rule generality. We fix $\tau_{feat}=10.0$ to limit the number of conjuncts per rule to at most ten, a choice motivated by default max-length settings in popular rule-mining libraries and grounded in human cognitive constraints defined by Miller’s Law [19] of $7\pm 2$ chunks for short-term memory.
@@
These conditions are applied sequentially. The overall objective function is the average over all $N$ instances: $\bar{M}=\frac{1}{N}\sum_{n=1}^{N}M(n).$ Maximizing $\bar{M}$ finds hyperparameters balancing COV(n){COV}(n) (not too low) and CONF(n){CONF}(n) (not too low), avoiding empty solutions, and minimizing rule complexity.

### v4

@@ 3.1 Overview of the PHAR Framework
The PHAR (Post-hoc Attribution Rules for time series) framework addresses the demand for semantically meaningful, stable, and cognitively efficient explanations in time series classification. In PHAR, we transform post-hoc, instance-level numeric feature attributions into concise, human-readable rule-based explanations and integrate outputs from multiple explainers enabling assessment of the quality of explanations and their visualization. First, PHAR selects the most salient temporal intervals—and, in multivariate settings, the most informative channels—so that rules focus on key patterns. Next, these intervals are translated into interpretable “IF–THEN” rules that preserve the model’s predicted class. An optional fusion step then merges rules obtained from numeric attributions, Anchor explanations [Ribeiro_Singh_Guestrin_2018], and potentially expert rules, optimizing for coverage, confidence, and simplicity. Fusing explanations via techniques such as intersection, union, and weighted aggregation ensures faithfulness to DL classifier predictions and usability for end-users. Finally, to reduce extraneous cognitive load, PHAR visualizes each rule with semi-factual plots that contrast actual and hypothetical behaviors, highlighting how small changes would affect the model’s decision. Our framework is structured around 4 Research Questions that guide the transformation, fusion, and visualization of rule-based explanations for time series classification.
@@
RQ3: How can fusion methods, combining rules from LIME, SHAP, and Anchor or from a single explainer, produce an optimized rule per instance that improves coverage, confidence, and explanation ratio while controlling rule complexity?
RQ4: How can visualizations support the interpretability of rule-based explanations in TS models? Rules may span multiple time steps and features (in multivariate TS). We explore whether overlaying rule-based explanations on TS plots helps to illustrate their structure.
Figure 2: Overview of the PHAR framework and its main stages (RQ1–RQ4). Numeric explainers (e.g. LIME, SHAP) and rule-based methods (e.g. Anchor) are shown as illustrative modules that can be replaced by other compatible explainers.
Figure 2 presents PHAR as a modular framework aligned with our Research Questions. A deep classifier first assigns each input time series, univariate or multivariate, to a predicted class. Local post-hoc explainers, such as SHAP and LIME, produce numeric attributions that we convert into interval rules of the form $r:\;x_{c}^{(t)}\in(l_{c}^{(t)},h_{c}^{(t)}]\;\Rightarrow\;\hat{y},\quad\text{conf}(r)=\gamma,\;\text{cov}(r)=\kappa,$ where $x_{c}^{(t)}$ denotes channel $c$ (for multivariate TS) at time step $t$, $(l_{c}^{(t)},h_{c}^{(t)}]$ is the selected interval and $\gamma,\kappa$ are the confidence and coverage of the rule. This transformation from numeric-to-rule addresses RQ1 (Section 3.3).
In parallel, a native rule-based explanation, such as Anchor, generates rules directly from the classifier. All candidate rules, regardless of origin, are then evaluated using metrics such as coverage, confidence, and feature count, which addresses RQ2 (Section 3.3.1).
Next, PHAR applies multiple fusion strategies (Intersection, Union, Weighted, Lasso, Lasso global, and Best) to these rule sets, using the same metrics to select a single optimized rule per instance (RQ3, Section 3.4).
Finally, to support RQ4, the resulting rules are visualized as vertical intervals overlayed on the original time series, highlighting which channels and timesteps support the prediction (Section 4.7 and Appendix A.4).
Although Figure 2 instantiates the framework with SHAP, LIME, and Anchor, the same pipeline can integrate other numeric or rule-based explanations as configurable elements.
@@ 3.2 Rule-based explanations
Converting scores from explainers that produce raw post hoc numerical attributions, such as SHAP and LIME, into a structured rule format makes it possible to compare and merge rules from different explainers, resolve conflicts when features are correlated, and engage Domain Experts in validating and refining the symbolic rules. Our approach employs rule fusion to aggregate rule sets produced by post hoc explainers and additionally allows us to incorporate native rule-based explainers like Anchor or even rules developed directly by Domain Experts. Inspired by ensemble methods [dietterich2000ensemble], we call this process fusion because it goes beyond classical bagging or boosting. Fusion refines and unifies rules by weighting them according to coverage, precision, and simplicity, and by merging redundant or overlapping conditions. The resulting fused rules conform to the same expert system format, allowing them to be easily evaluated, adjusted, visualized, and aligned with expert knowledge to support robust domain reasoning.
In Davis1977, rules are expressed in the form "IF–THEN" augmented with a certainty factor. The premise is a conjunction of predicate evaluations $p_{j}(o_{j},a_{j})=v_{j}$ for $j=1,\dots,n$, and the conclusion $q(o^{\prime},a^{\prime})=v^{\prime}_{\mathrm{out}}$ is drawn with a strength given by multiplying the aggregated premise confidence by $\mathrm{CF}_{i}$:
@@
Here, $p_{j}$ denotes a predicate function applied to object $o_{j}$ and attribute $a_{j}$, $v_{j}$ is its value, and the certainty factor $\mathrm{CF}_{i}$ quantifies the reliability of the inference. Building on this idea, we define our PHAR rule format as in [ligeza_rule_based, Niederlinski2013], where each rule $R_{i}$ comprises a set of logical conditions $W_{1},\dots,W_{n}$, a conclusion $W_{\mathrm{out}}$, and an inference semaphore $S_{i}$ representing the combined confidence $c_{i}$ obtained with the method described in Section 3.3.1:
@@
In this schema, each $W_{j}$ is a logical condition derived from feature attributions, $W_{\mathrm{out}}$ is the predicted class or decision, and $S_{i}=c_{i}$ signals the confidence of the rule.
Our approach follows established rule classifications in knowledge-based systems [Nalepa2018]. The PHAR rules are best described as derivation and classification rules. Derivation rules define logical conditions under which a conclusion is inferred from given premises, in our case TS feature conditions leading to model predictions. Classification rules assign instances to classes based on these conditions. From the perspective of logical inference, our rules are probabilistic deductive rules. Deductive rules infer conclusions from premises. If the premises hold, the conclusion follows deductively. Given a set of features and its constraints, our rules classify instances under a probabilistic confidence threshold rather than absolute certainty. Unlike classical deduction, our rules incorporate confidence, applying deterministically within probability constraints. Furthermore, our rules belong to attributive rule languages [Nalepa2018]. These languages encode knowledge through attributes and values, which are common in decision tables, rule-based systems, and databases. They enable intuitive representation and efficient manipulation. For example, the rule if $\mathrm{temperature}(t=10)>50$ and $\mathrm{pressure}(t=11)<10$, then class = failure [$c=0.85$], where $t=11$ denotes the timestep index, provides a human-readable decision boundary that replaces raw numeric attribution scores.
@@
Here, each pair $t_{j_{m}}c_{j_{m}}$ denotes “channel” $c_{j_{m}}$ (in multivariate series) at the time step $t_{j_{m}}$. The indices $j_{1},\dots,j_{k}$ are arbitrarily selected from the complete set of $T$ time steps and $C$ "channels", choosing the pairs $k+1$ with the greatest impact on model predictions. Each interval $(\ell_{j_{m}},u_{j_{m}}]$ has the lower bound $\ell_{j_{m}}$ and the upper bound $u_{j_{m}}$. $\mathrm{CONF}(R_{i})$ is the proportion of instances that satisfy the rule that are correctly classified as class $A"$, and $\mathrm{COV}(R_{i})$ is the fraction of all instances that satisfy the rule. The formal derivation of the rules and their CONF and COV measures is presented in Sections 3.3 and 3.3.1, while Section 4.7 provides real-world examples of the rules in this notation for a selected dataset.
@@ 3.3 Conversion of numeric importance values to rules
We consider a TS dataset with $N$ samples and a DL model that produces multi-class predictions. In the univariate case, each feature $f$ is indexed only by timestep $\mathrm{t}_{i}$, whereas in the multivariate case, the features are indexed by timestep and channel, for example $\mathrm{t}_{i}\mathrm{c}_{v}$. Let $\{X_{n}\}_{n=1}^{N}$ denote the input instances with model predictions $y_{n}$. In this work, we explain the behavior of the trained model, so the predicted classes $y_{n}$ act as a ground truth for the explanation stage.
Feature-based explainers, such as LIME and SHAP, assign a numeric importance value to each feature of each instance. Let $e_{n,f}$ denote the importance of the feature $f$, for instance $n$. To extract rules, we first identify a threshold $T$ that marks important features based on $|e_{n,f}|$. In PHAR, $T$ is percentile-based and can be computed per feature (local threshold) or across all features (global threshold). Other thresholding strategies are also possible, for example, selecting the top-$k$ features per instance, using a fixed absolute cutoff on $|e_{n,f}|$, or defining $T$ from distributional statistics such as $\mu+\lambda\sigma$. However, in this study, we do not explore these alternatives.
We first compute absolute explanation values for all training instances and features. Let $E$ be the multiset of absolute values: $E=\{|e_{n,f}|:n=1,\dots,N,\;f=1,\dots,F\},$ where $e_{n,f}$ is the explanation value for the feature $f$ (e.g. $\text{t}_{i}\text{c}_{v}$ for multivariate or $\text{t}_{i}$ for univariate TS) at instance $n$. We then define a global threshold $T_{global}$ by selecting a percentile $p$: $T_{global}=\text{Percentile}(E,p).$ If we apply thresholds per feature, we instead compute a separate threshold $T_{f}$ for each feature $f$: $E_{f}=\{|e_{n,f}|:n=1,\dots,N\},\quad T_{f}=\text{Percentile}(E_{f},p).$ During inference in test instances, for each feature $f$, let $e_{n,f}$ be the explanation value for instance $n$:
@@
Features that exceed their threshold are important. For important features, we generate intervals that preserve the prediction of the model. Stable intervals $(l_{f},u_{f}]$ are derived by controlled perturbation of each important feature and sampling around $x_{n,f}$. Let $F_{*}$ be important features for an instance, identified after thresholding. For each $f\in F_{*}$, we take the original value $x_{n,f}$ and generate perturbed samples: $x_{n,f}^{(m)}\sim\text{Uniform}(x_{n,f}-\Delta_{f},\;x_{n,f}+\Delta_{f}),$ where $\Delta_{f}$ is a perturbation range based on training explanation statistics (std. deviation × constant $\sigma_{p}$). Joint sampling over $f\in F_{*}$ captures nonlinear feature dependencies. For each perturbed sample $(x_{n,f}^{(m)})_{f\in F_{*}}$, we calculate the prediction of the model. Let $c_{ref}(n)$ be the original predicted class of instance $n$. If $c_{ref}(n)=y_{n}$, we determine $(l_{f},u_{f}]$ per feature $f$ that preserves this class. Each $f\in F_{*}$ receives a stable interval: $f:x_{f}\in(l_{f},u_{f}],$ which defines a rule $R_{n}$ for the instance $n$:
@@ 3.3.1 Confidence and coverage of newly derived rules
After deriving the final rule $R_{n}$ for the instance $n$, we evaluate its explanatory quality on the test set. Consider the test data set $\{X_{test,k}\}_{k=1}^{N_{test}}$, where $N_{test}$ is the number of test instances. We identify the subset of test instances that satisfy the rule: $S(R_{n})=\{X_{test,k}:X_{test,k}\text{ satisfies }R_{n}\}.$ To measure coverage, we compute the following:
@@
This ratio indicates how many test instances fall into the region defined by the rule’s intervals. A higher $COV(R_{n})$ means that the rule covers a larger portion of the test set.
To measure confidence, we focus on the consistency of the predictions within $S(R_{n})$. Let: $T(R_{n})=\{X_{test,k}\in S(R_{n}):y_{model}(X_{test,k})=c_{ref}(n)\},$ i.e., the subset of test instances covered for which the predicted class by the model matches $c_{ref}(n)$. The confidence is then:
@@
If no test instance satisfies the rule ($S(R_{n})=\emptyset$), then $COV(R_{n})=0$ and $CONF(R_{n})$ are undefined, which means that there is no explanatory power. However, at least the initial instance $n$ should satisfy the rule. If the rule covers test instances but does not match $c_{ref}(n)$, confidence decreases.
@@ 3.3.2 Hyperparameter optimization in rule transformation step
To ensure that our rule‐-based explanations achieve a balance between fidelity and simplicity, we employ Optuna  [^2] Optuna: https://optuna.org/ [optuna_2019] to tune key hyperparameters of this step of our framework.
@@
  • $p$ (threshold percentile): Feature importance threshold in terms of percentile. Higher values prioritize top-scoring features and thus result in a lower number of important features.
@@
  • $\sigma_{p}$ (perturbation scale): Perturbation scale. Higher values expand the search for stable intervals.
  • $N_{p}$ (count of perturbation samples): Number of perturbed samples for interval estimation. More samples improve precision, but increase computational cost.
Optuna searches over the hyperparameters $p$, $\delta$, $\sigma_{p}$ and $N_{p}$ to maximize the objective in a holding ‐out test set. During each trial, candidate configurations are evaluated by applying uniform sampling to estimate rule stability and decision region consistency, ensuring stable decision regions in the input space of the ML model. The best configuration is selected based on its performance in the test set, ensuring that the chosen thresholds generalize beyond the training data. The tuning of these parameters balances feature selectivity, interval granularity, and computational cost. Table 1 (Appendix) summarizes the hyperparameter ranges used in our study.
Specifically, we define an objective function maximizing coverage and confidence while penalizing empty or complex rules during numeric explanations to the rule transformation step which is then fed to Optuna framework.
@@
Each instance $n$ starts with a complex rule from an explainer (e.g. Anchor, LIME, SHAP). Each rule defines intervals for features, e.g. timestep-channel $\text{t}_{i}\text{c}_{v}$ in multivariate TS or timestep $\text{t}_{i}$ in univariate TS, denoted $F(R_{n})$. The rules yield metrics: coverage $\operatorname{COV}(R_{n})$, confidence $\operatorname{CONF}(R_{n})$, and count of features that the rule specifies $|F(R_{n})|$. The initial objective function is as follows:
@@
If rule for the instance $n$ is empty or absent, $M(n)=0$. Otherwise, penalties are applied when certain thresholds are not met.
Then we introduce the thresholds $\tau_{conf}$, $\tau_{cov}$, and $\tau_{feat}$, which are chosen to balance the quality and interpretability of the rules. $\tau_{conf}$ ensures a minimum confidence level for the rules, $\tau_{cov}$ prevents overly narrow rules by enforcing minimum coverage, and $\tau_{feat}$ limits the number of features in a rule to improve cognitive interpretability for end-users. We set $\tau_{conf}=0.5$ because scikit-learn and many binary classifiers use 0.5 as the default probability cutoff, aligning the applicability of the rule with positive predictions above the chance level and following the calibration practices. The coverage threshold $\tau_{cov}=0.01$ (1% support) corresponds to the minimum support filter commonly adopted in association rule mining to remove infrequent patterns and ensure rule generality. We fix $\tau_{feat}=10.0$ to limit the number of conjuncts per rule to at most ten, a choice motivated by default max-length settings in popular rule-mining libraries and grounded in human cognitive constraints defined by Miller’s Law [Miller1956] of $7\pm 2$ chunks for short-term memory.
@@
These conditions are applied sequentially. The overall objective function is the average over all $N$ instances: $\bar{M}=\frac{1}{N}\sum_{n=1}^{N}M(n).$ Maximizing $\bar{M}$ finds hyperparameters balancing COV(n){COV}(n) (not too low) and CONF(n){CONF}(n) (not too low), avoiding empty solutions and minimizing the complexity of the rules.